OpenStudy (anonymous):

Solve the system x - y = 3 and 3x - 8y = 9. What is the value of y?

OpenStudy (anonymous):

My last question guys! Promise!!

OpenStudy (anonymous):

x = y + 3 plug this in for x into the other equation

hero (hero):

x - y = 3 3x - 8y = 9 x -3 = y 3x - 9 = 8y y = x - 3 y = (3x-9)/8 y = y 8(x-3) = 3x-9 8x - 24 = 3x - 9 8x - 3x = 24 - 9 5x = 15 x = 15/5 x = 3 y = 0

hero (hero):

(x,y) = (3,0)